# Ctrl-GenAug: Controllable Generative Augmentation for Medical Sequence Classification

### v2

@@ 4.1 Datasets and Settings
Method Backbone Training Paradigm Carotid TUSC (Stanford-AIMI, 2021) ACDC (Bernard et al., 2018)
I3D  (Carreira and Zisserman, 2017) CNN A 79.59 0.737 60.72 0.559 83.49 0.733
B 85.03 $\uparrow$ 5.44 0.813 $\uparrow$ 0.076 65.41 $\uparrow$ 4.69 0.611 $\uparrow$ 0.052 87.16 $\uparrow$ 3.67 0.795 $\uparrow$ 0.062
C 84.35 $\uparrow$ 4.76 0.811 $\uparrow$ 0.074 62.88 $\uparrow$ 2.16 0.578 $\uparrow$ 0.019 85.32 $\uparrow$ 1.83 0.762 $\uparrow$ 0.029
R(2+1)D  (Tran et al., 2018) CNN A 82.99 0.781 63.96 0.595 83.85 0.737
B 83.67 $\uparrow$ 0.68 0.797 $\uparrow$ 0.016 64.68 $\uparrow$ 0.72 0.601 $\uparrow$ 0.006 85.69 $\uparrow$ 1.84 0.769 $\uparrow$ 0.032
C 85.04 $\uparrow$ 2.05 0.804 $\uparrow$ 0.023 65.41 $\uparrow$ 1.45 0.610 $\uparrow$ 0.015 85.69 $\uparrow$ 1.84 0.771 $\uparrow$ 0.034
SlowFast  (Feichtenhofer et al., 2019) CNN A 80.95 0.742 62.88 0.579 81.65 0.702
B 81.63 $\uparrow$ 0.68 0.770 $\uparrow$ 0.028 65.76 $\uparrow$ 2.88 0.609 $\uparrow$ 0.030 86.05 $\uparrow$ 4.40 0.772 $\uparrow$ 0.070
C 84.35 $\uparrow$ 3.40 0.811 $\uparrow$ 0.069 65.41 $\uparrow$ 2.53 0.609 $\uparrow$ 0.030 83.85 $\uparrow$ 2.20 0.742 $\uparrow$ 0.040
CSN  (Tran et al., 2019) CNN A 80.27 0.757 61.80 0.570 85.69 0.765
B 84.35 $\uparrow$ 4.08 0.784 $\uparrow$ 0.027 67.21 $\uparrow$ 5.41 0.631 $\uparrow$ 0.061 87.15 $\uparrow$ 1.46 0.794 $\uparrow$ 0.029
C 87.07 $\uparrow$ 6.80 0.843 $\uparrow$ 0.086 65.41 $\uparrow$ 3.61 0.610 $\uparrow$ 0.040 88.62 $\uparrow$ 2.93 0.814 $\uparrow$ 0.049
TPN  (Yang et al., 2020) CNN A 82.99 0.784 64.32 0.602 83.85 0.735
B 87.76 $\uparrow$ 4.77 0.856 $\uparrow$ 0.072 66.85 $\uparrow$ 2.53 0.633 $\uparrow$ 0.031 85.32 $\uparrow$ 1.47 0.764 $\uparrow$ 0.029
C 85.71 $\uparrow$ 2.72 0.823 $\uparrow$ 0.039 65.41 $\uparrow$ 1.09 0.609 $\uparrow$ 0.007 84.95 $\uparrow$ 1.10 0.756 $\uparrow$ 0.021
TimeSformer  (Bertasius et al., 2021) Transformer A 77.55 0.721 65.41 0.614 72.48 0.562
B 80.95 $\uparrow$ 3.40 0.753 $\uparrow$ 0.032 66.49 $\uparrow$ 1.08 0.626 $\uparrow$ 0.012 76.51 $\uparrow$ 4.03 0.626 $\uparrow$ 0.064
C 79.59 $\uparrow$ 2.04 0.730 $\uparrow$ 0.009 66.49 $\uparrow$ 1.08 0.618 $\uparrow$ 0.004 77.62 $\uparrow$ 5.14 0.642 $\uparrow$ 0.080
MViTv2  (Li et al., 2022b) Transformer A 74.15 0.663 59.64 0.543 77.98 0.642
B 78.91 $\uparrow$ 4.76 0.740 $\uparrow$ 0.077 63.96 $\uparrow$ 4.32 0.596 $\uparrow$ 0.053 78.72 $\uparrow$ 0.74 0.653 $\uparrow$ 0.011
C 76.87 $\uparrow$ 2.72 0.701 $\uparrow$ 0.038 62.17 $\uparrow$ 2.53 0.568 $\uparrow$ 0.025 80.18 $\uparrow$ 2.20 0.680 $\uparrow$ 0.038
VideoSwin  (Liu et al., 2022) Transformer A 79.59 0.732 58.92 0.535 74.31 0.572
B 83.67 $\uparrow$ 4.08 0.789 $\uparrow$ 0.057 61.08 $\uparrow$ 2.16 0.551 $\uparrow$ 0.016 79.45 $\uparrow$ 5.14 0.663 $\uparrow$ 0.091
C 83.67 $\uparrow$ 4.08 0.776 $\uparrow$ 0.044 61.44 $\uparrow$ 2.52 0.562 $\uparrow$ 0.027 77.62 $\uparrow$ 3.31 0.632 $\uparrow$ 0.060
UniFormerV2  (Li et al., 2022a) Hybrid A 76.87 0.710 65.41 0.611 75.78 0.614
B 80.27 $\uparrow$ 3.40 0.743 $\uparrow$ 0.033 65.77 $\uparrow$ 0.36 0.618 $\uparrow$ 0.007 77.25 $\uparrow$ 1.47 0.645 $\uparrow$ 0.031
C 81.63 $\uparrow$ 4.76 0.760 $\uparrow$ 0.050 66.13 $\uparrow$ 0.72 0.617 $\uparrow$ 0.006 77.61 $\uparrow$ 1.83 0.637 $\uparrow$ 0.023
FTC  (Ahmadi et al., 2023) Hybrid A 76.99 0.723 62.09 0.584 80.95 0.689
B 80.91 $\uparrow$ 3.92 0.750 $\uparrow$ 0.027 64.09 $\uparrow$ 2.00 0.593 $\uparrow$ 0.009 82.38 $\uparrow$ 1.43 0.703 $\uparrow$ 0.014
C 81.80 $\uparrow$ 4.81 0.758 $\uparrow$ 0.035 65.66 $\uparrow$ 3.57 0.614 $\uparrow$ 0.030 84.38 $\uparrow$ 3.43 0.718 $\uparrow$ 0.029
CSG-3DCT  (Zhou et al., 2023) Hybrid A 84.35 0.812 61.44 0.571 87.16 0.797
B 87.07 $\uparrow$ 2.72 0.831 $\uparrow$ 0.019 63.96 $\uparrow$ 2.52 0.590 $\uparrow$ 0.019 88.62 $\uparrow$ 1.46 0.819 $\uparrow$ 0.022
C 87.76 $\uparrow$ 3.41 0.837 $\uparrow$ 0.025 65.05 $\uparrow$ 3.61 0.609 $\uparrow$ 0.038 88.63 $\uparrow$ 1.47 0.820 $\uparrow$ 0.023
4.2 Generative Augmentation-based Sequence Classification
Table 2 compares the average performance (i.e., accuracy and AUROC) on different training paradigms across a wide range of sequence classifiers, including five CNN-based classic models, three transformer-based models, and three CNN-Transformer hybrid-based methods. It is evident that the performance ceiling of existing popular models, from simple CNN-based to complex CNN-Transformer hybrids, can be further elevated using synthetic samples obtained from our proposed framework, under Real-finetune and Joint-train paradigms. On the carotid dataset, the accuracy/AUROC can be most boosted by 5.44%/0.077 and 6.80%/0.086 through Real-finetune and Joint-train paradigms, respectively. Due to high intra-class variations in TUSC, the baseline exhibits visibly inferior results, with an average accuracy of 62.42% compared to other datasets. However, it can be seen from Table 2 that samples synthesized by Ctrl-GenAug make significant contributions and enhance the classification performance to an acceptance level. Among all the classifiers, CSN harvests the largest performance gain, which is non-trivially enhanced by 5.41%/0.061 and 3.61%/0.040 accuracy/AUROC under Real-finetune and Joint-train paradigms, respectively. Notably, for ACDC, the overall performance improvement remains significant even with a high baseline. For instance, CSG-3DCT is boosted from 87.16$\rightarrow$88.63(+1.47%) accuracy and from 0.797$\rightarrow$0.820 (+0.023) AUROC under the Joint-train practice. Extensive experimental results show that synthetic sample sets produced by Ctrl-GenAug can be powerful data engines to enhance various diagnostic tasks in medical sequence analysis.
@@ 4.3 Improved Diagnosis in Underrepresented High-risk Sets
Fig. 7 quantitatively compares the carotid stenosis diagnostic performance of underrepresented high-risk sets on Baseline and Joint-train paradigms. It shows that Baseline performs poorly on underrepresented high-risk sets, especially in terms of sensitivity (average 18.52% in moderate and 36.25% in severe sets). In comparison, the classifiers can be greatly enhanced by jointly training with synthetic samples (e.g., SlowFast: 11.11$\rightarrow$44.44sensitivity on the moderate set).
We also provided the category-level analysis, and summarized the average F1 scores across four classifiers for a global observation and conclusion. Specifically, Baseline reached 69.92%, 25.92%, and 43.44% average F1 scores for mild, moderate, and severe cases, respectively. The F1 performance of Joint-train in these three categories was 82.44%, 52.73%, and 60.40%. We found that well-designed classifiers jointly trained on synthetic samples from Ctrl-GenAug and real ones, can effectively close diagnostic performance gaps between overrepresented (i.e., mild) and underrepresented sets (i.e., moderate and severe) while improving the former ($\sim$13$\uparrow$).
Figure 7: Carotid performance radar map, using 4 models trained on Baseline (green) and Joint-train (blue) paradigms in underrepresented high-risk sets.
@@ 4.4 Classification Robustness in Out-domain Conditions
We further explored the impact of synthetic samples from Ctrl-GenAug on diagnostic robustness in out-domain settings using the carotid dataset. Training data of the sequence generator consisted of labeled ID and unlabeled OD subsets. We added the hospital identifier in the descriptive text of the ID data to condition the domain distribution. For the OD data, which did not contain the diagnostic class label, we padded the corresponding conditioning vector with zeros, while solely preserving hospital IDs to form descriptive texts. As shown in Table 3, the average accuracy and AUROC of three classifiers display better results on the Real-finetune and Joint-train paradigms compared to the Baseline. It validates that diagnostic robustness can be enhanced with the aid of Ctrl-GenAug in scenarios where we only have access to unlabeled cases from additional medical centers due to limited resources.
@@
Method Backbone Training Paradigm Carotid
Acc. $\Delta$ AUROC $\Delta$
I3D  (Carreira and Zisserman, 2017) CNN A 70.00 0.609
B 71.67 $\uparrow$ 1.67 0.648 $\uparrow$ 0.039
C 75.00 $\uparrow$ 5.00 0.682 $\uparrow$ 0.073
TPN  (Yang et al., 2020) CNN A 75.00 0.667
B 76.67 $\uparrow$ 1.67 0.731 $\uparrow$ 0.064
C 76.67 $\uparrow$ 1.67 0.678 $\uparrow$ 0.011
CSG-3DCT  (Zhou et al., 2023) Hybrid A 68.33 0.599
B 73.33 $\uparrow$ 5.00 0.689 $\uparrow$ 0.090
C 70.00 $\uparrow$ 1.67 0.651 $\uparrow$ 0.052
Table 4: Ablation study for different conditional controls in TUSC (Stanford-AIMI, 2021), including class label (C), text (T), image prior (I), and motion field (MF). SlowFast (Feichtenhofer et al., 2019) (C1), CSN (Tran et al., 2019) (C2), and FTC (Ahmadi et al., 2023) (C3) trained on Joint-train paradigm were used for downstream diagnosis. “Hybrid”, the classifier with CNN-Transformer design.
Dataset Synthesis Downstream Diagnosis
Control Metric Backbone $\Delta$Accuracy $\Delta$AUROC
[…]
@@
4.5 Analysis of Multimodal Conditions Guidance
To illustrate the role of different conditional guidance, we conduct comparative experiments with generators that resort to various banks of conditions for training and sampling. Table 4 validates that all proposed conditions are effective for enhancing synthesis and diagnosis tasks. Specifically, without text control, the downstream diagnostic performance is comparatively poor. Fig. 6(b-c) compares typical synthetic thyroid nodule sequences generated under condition banks with and without text guidance. The former is visibly high-fidelity and faithful to the given text prompt (e.g., smooth margin), while the latter exhibits less distinguishable features for diagnosis (e.g., blurry margin). That is to say, the proposed Ctrl-GenAug can create more diagnosis-reliable samples by enhancing semantic steerability in the generation process. Besides, by incorporating image prior knowledge in the sampling process, the average accuracy and AUROC of three classifiers are improved by 4.20% and 0.040, respectively. This proves the domain gap between synthetic and real samples is mitigated by introducing the image prior. Moreover, as shown in Table 4, conditioning the generator on the motion field produces more content-consistent and smoother samples, leading to continuous improvements in diagnostic performance. It can also be observed that the FVD results show a limited correlation with downstream evaluation metrics, confirming the finding in (Luo et al., 2024).
Table 5: Ablation study for the sequential augmentation module of the generator in TUSC (Stanford-AIMI, 2021) and ACDC (Bernard et al., 2018). SlowFast (Feichtenhofer et al., 2019) (C1), CSN (Tran et al., 2019) (C2), and FTC (Ahmadi et al., 2023) (C3) trained on Joint-train paradigm were used for downstream diagnosis.
“Hybrid”, the classifier with CNN-Transformer design.
@@ 4.5 Analysis of Multimodal Conditions Guidance
4.6 Impact of the Sequential Augmentation Module
To verify the impact of the proposed sequential augmentation module, as shown in Table 5, we compared the performance of three generator variants on the synthesis and downstream tasks. Ours-S, Ours-SK, and Ours-SKM denote our ablation studies, including gradually adding sequential attention (’-S’), key-frame/slice attention (’-K’), and motion field attention (’-M’) to the generator with spatial inflation only. Table 5 shows that the downstream accuracy and AUROC improvements of Ours-SKM are significantly higher than those of Ours-S and Ours-SK in both datasets. In terms of quantitative assessment of sequential coherence, Ours-SKM achieves better results overall than other variants. These prove that this module can help the generator synthesize cross-frame/slice consistent and dynamic-smooth clips, which are diagnosis-promotive for downstream diagnostic tasks. Moreover, it can be observed that the FVD results across these three generators show no significant difference, revealing its limited value for evaluating diagnosis-oriented medical sequence synthesis tasks.
4.7 Effectiveness of the Noisy Synthetic Data Filter

### v3

@@ 4.5 Generative Augmentation for Diagnosis with Multiple Sequence Classifiers
Table 5: Out-domain performance comparison using eleven nets trained on Baseline (A), Real-finetune (B), and Joint-train (C) paradigms in carotid dataset. Underlined results show no significant difference compared with Baseline (A), all others indicate significant differences ($p<0.05$). 95% confidence intervals for both metrics are reported. “Hybrid”, the net with CNN-Transformer design. Acc., accuracy ($\%$).
Classifier Backbone Training Paradigm Carotid
Acc. $\Delta$ AUROC $\Delta$
I3D (Carreira and Zisserman, 2017) CNN A 70.00 [60.00, 80.00] 0.609 [0.509, 0.721]
B 71.67 [61.67, 81.67] $\uparrow$ 1.67 0.648 [0.537, 0.771] $\uparrow$ 0.039
C 75.00 [65.00, 85.00] $\uparrow$ 5.00 0.682 [0.565, 0.794] $\uparrow$ 0.073
R(2+1)D (Tran et al., 2018) CNN A 66.67 [56.67, 76.71] 0.577 [0.471, 0.690]
B 71.67 [61.67, 81.67] $\uparrow$ 5.00 0.661 [0.556, 0.773] $\uparrow$ 0.084
C 75.00 [65.00, 85.00] $\uparrow$ 8.33 0.689 [0.575, 0.810] $\uparrow$ 0.112
SlowFast (Feichtenhofer et al., 2019) CNN A 73.33 [63.33, 83.33] 0.638 [0.528, 0.750]
B 75.00 [65.00, 85.00] $\uparrow$ 1.67 0.670 [0.562, 0.787] $\uparrow$ 0.032
C 75.00 [65.00, 85.00] $\uparrow$ 1.67 0.706 [0.588, 0.817] $\uparrow$ 0.068
CSN (Tran et al., 2019) CNN A 68.33 [58.33, 78.33] 0.592 [0.494, 0.703]
B 73.33 [63.33, 83.33] $\uparrow$ 5.00 0.664 [0.556, 0.786] $\uparrow$ 0.072
C 71.67 [61.67, 81.67] $\uparrow$ 3.34 0.672 [0.552, 0.784] $\uparrow$ 0.080
TPN (Yang et al., 2020) CNN A 75.00 [65.00, 85.00] 0.667 [0.563, 0.783]
B 76.67 [66.67, 86.67] $\uparrow$ 1.67 0.731 [0.624, 0.831] $\uparrow$ 0.064
C 76.67 [66.67, 86.67] $\uparrow$ 1.67 0.678 [0.569, 0.790] $\uparrow$ 0.011
TimeSformer (Bertasius et al., 2021) Transformer A 68.33 [58.33, 78.33] 0.590 [0.489, 0.689]
B 71.67 [61.62, 81.67] $\uparrow$ 3.34 0.665 [0.543, 0.787] $\uparrow$ 0.075
C 73.33 [63.33, 83.33] $\uparrow$ 5.00 0.664 [0.565, 0.773] $\uparrow$ 0.074
MViTv2 (Li et al., 2022b) Transformer A 66.67 [56.67, 76.67] 0.590 [0.480, 0.699]
B 68.33 [58.33, 78.33] $\uparrow$ 1.66 0.592 [0.493, 0.699] $\uparrow$ 0.002
C 70.00 [60.00, 80.00] $\uparrow$ 3.33 0.604 [0.508, 0.706] $\uparrow$ 0.014
VideoSwin (Liu et al., 2022) Transformer A 65.00 [55.00, 75.00] 0.574 [0.462, 0.688]
B 70.00 [58.33, 80.00] $\uparrow$ 5.00 0.650 [0.527, 0.767] $\uparrow$ 0.076
C 68.33 [58.33, 78.33] $\uparrow$ 3.33 0.609 [0.503, 0.728] $\uparrow$ 0.035
UniFormerV2 (Li et al., 2022a) Hybrid A 66.67 [56.67, 76.67] 0.617 [0.498, 0.735]
B 73.33 [63.33, 83.33] $\uparrow$ 6.66 0.681 [0.564, 0.803] $\uparrow$ 0.064
C 68.33 [58.33, 78.33] $\uparrow$ 1.66 0.633 [0.521, 0.756] $\uparrow$ 0.016
FTC (Ahmadi et al., 2023) Hybrid A 67.08 [56.67, 78.33] 0.616 [0.502, 0.741]
B 71.31 [66.67, 80.00] $\uparrow$ 4.23 0.632 [0.523, 0.752] $\uparrow$ 0.016
C 74.68 [66.67, 83.33] $\uparrow$ 7.60 0.670 [0.553, 0.784] $\uparrow$ 0.054
CSG-3DCT (Zhou et al., 2023) Hybrid A 68.33 [58.33, 78.33] 0.599 [0.491, 0.715]
B 73.33 [63.33, 83.33] $\uparrow$ 5.00 0.689 [0.569, 0.805] $\uparrow$ 0.090
C 70.00 [60.00, 80.00] $\uparrow$ 1.67 0.651 [0.530, 0.772] $\uparrow$ 0.052
@@ 4.6 Improved Diagnosis in Underrepresented High-risk Sets
[…]
@@ 4.7 Classification Robustness in Out-domain Conditions
We further explored the impact of synthetic samples from Ctrl-GenAug on diagnostic robustness in out-domain settings using the carotid dataset. Training data of the sequence generator consisted of labeled ID and unlabeled OD subsets. We added the hospital identifier in the descriptive text of the ID data to condition the domain distribution. For the OD data, which did not contain the diagnostic class label, we padded the corresponding conditioning vector with zeros, while solely preserving hospital IDs to form descriptive texts. As shown in Table 5, the average accuracy and AUROC of all eleven classifiers display better results on the Real-finetune and Joint-train paradigms compared to the Baseline, with most of the improvements showing statistical significance ($p<0.05$). It validates that diagnostic robustness can be enhanced with the aid of Ctrl-GenAug in scenarios where we only have access to unlabeled cases from additional medical centers due to limited resources.
Table 6: Ablation study for different conditional controls in multi-organ and multi-modal datasets, including class label (C), text (T), image prior (I), and motion field (MF). SlowFast (Feichtenhofer et al., 2019) (C1), CSN (Tran et al., 2019) (C2), and FTC (Ahmadi et al., 2023) (C3) trained on Joint-train paradigm were used for downstream diagnosis. The accuracy and AUROC gains over the Baseline are reported. “Hybrid”, the classifier with CNN-Transformer design.
Dataset Synthesis Downstream Diagnosis
Control Metric Backbone $\DeltaAccuracy Δ\DeltaAUROC$
@@
Carotid 4.18 89.65% 93.27% C1 CNN $\uparrow$ 2.04% $\uparrow$ 0.009
C2 CNN $\downarrow$ 0.68% $\downarrow$ 0.009
C3 Hybrid $\uparrow$ 0.56% $\uparrow$ 0.031
6.50 89.08% 92.85% C1 CNN $\uparrow$ 1.36% $\uparrow$ 0.046
C2 CNN $\downarrow$ 0.68% $\downarrow$ 0.011
C3 Hybrid $\uparrow$ 1.12% $\uparrow$ 0.021
4.29 88.74% 93.09% C1 CNN $\uparrow$ 2.72% $\uparrow$ 0.078
C2 CNN $\uparrow$ 4.09% $\uparrow$ 0.094
C3 Hybrid $\uparrow$ 3.28% $\uparrow$ 0.064
4.26 91.46% 94.29% C1 CNN $\uparrow$ 3.40% $\uparrow$ 0.069
C2 CNN $\uparrow$ 6.80% $\uparrow$ 0.086
C3 Hybrid $\uparrow$ 4.81% $\uparrow$ 0.035
@@
ACDC (Bernard et al., 2018) - 8.70 76.74% 92.50% C1 CNN $\uparrow$ 1.46% $\uparrow$ 0.010
C2 CNN $\uparrow$ 0.37% $\uparrow$ 0.027
C3 Hybrid $\uparrow$ 1.10% $\uparrow$ 0.051
- 10.81 79.66% 93.89% C1 CNN $\uparrow$ 0.37% $\uparrow$ 0.020
C2 CNN $\uparrow$ 0.74% $\uparrow$ 0.012
C3 Hybrid $\uparrow$ 1.83% $\uparrow$ 0.029
- 8.18 77.08% 92.77% C1 CNN $\uparrow$ 2.20% $\uparrow$ 0.040
C2 CNN $\uparrow$ 2.93% $\uparrow$ 0.049
C3 Hybrid $\uparrow$ 3.43% $\uparrow$ 0.029
MosMedData (Vladzymyrskyy et al., 2020) - 2.98 83.06% 95.27% C1 CNN $\uparrow$ 2.96% $\uparrow$ 0.004
C2 CNN $\uparrow$ 1.41% $\downarrow$ 0.091
C3 Hybrid $\uparrow$ 1.15% $\uparrow$ 0.070
- 5.05 84.72% 96.35% C1 CNN $\uparrow$ 1.93% $\downarrow$ 0.026
C2 CNN $\uparrow$ 0.38% $\downarrow$ 0.114
C3 Hybrid $\uparrow$ 1.81% $\uparrow$ 0.054
- 2.91 84.07% 96.33% C1 CNN $\uparrow$ 3.73% $\uparrow$ 0.028
C2 CNN $\uparrow$ 2.57% $\uparrow$ 0.005
C3 Hybrid $\uparrow$ 2.70% $\uparrow$ 0.044
MRNet (Bien et al., 2018) - 8.48 75.48% 92.90% C1 CNN $\uparrow$ 2.12% $\uparrow$ 0.018
C2 CNN $\uparrow$ 0.53% $\uparrow$ 0.045
C3 Hybrid $\uparrow$ 3.17% $\uparrow$ 0.057
- 11.67 76.67% 93.61% C1 CNN $\uparrow$ 1.59% $\uparrow$ 0.017
C2 CNN $\downarrow$ 2.12% $\downarrow$ 0.020
C3 Hybrid $\uparrow$ 1.06% $\uparrow$ 0.026
- 8.14 77.10% 93.81% C1 CNN $\uparrow$ 2.12% $\uparrow$ 0.030
C2 CNN $\uparrow$ 1.58% $\uparrow$ 0.009
C3 Hybrid $\uparrow$ 4.76% $\uparrow$ 0.035
Table 7: Ablation study for the sequential augmentation module of the generator in multi-organ and multi-modal datasets. SlowFast (Feichtenhofer et al., 2019) (C1), CSN (Tran et al., 2019) (C2), and FTC (Ahmadi et al., 2023) (C3) trained on Joint-train paradigm were used for downstream diagnosis. The accuracy and AUROC gains over the Baseline are reported. “Hybrid”, the classifier with CNN-Transformer design.
@@
Carotid Ours-S 4.53 85.39% 87.45% C1 CNN $\uparrow$ 1.40% $\uparrow$ 0.031
C2 CNN $\uparrow$ 1.12% $\uparrow$ 0.009
C3 Hybrid $\downarrow$ 0.68% $\downarrow$ 0.017
Ours-SK 4.75 89.03% 93.38% C1 CNN $\uparrow$ 2.04% $\uparrow$ 0.044
C2 CNN $\uparrow$ 1.36% $\uparrow$ 0.053
C3 Hybrid $\uparrow$ 2.72% $\uparrow$ 0.062
Ours-SKM 4.26 91.46% 94.29% C1 CNN $\uparrow$ 3.40% $\uparrow$ 0.069
C2 CNN $\uparrow$ 6.80% $\uparrow$ 0.086
C3 Hybrid $\uparrow$ 4.81% $\uparrow$ 0.035
@@
MosMedData (Vladzymyrskyy et al., 2020) Ours-S 5.17 75.23% 89.05% C1 CNN $\uparrow$ 1.54% $\downarrow$ 0.002
C2 CNN $\downarrow$ 0.13% $\downarrow$ 0.102
C3 Hybrid $\downarrow$ 0.76% $\downarrow$ 0.013
Ours-SK 2.98 84.19% 96.16% C1 CNN $\uparrow$ 2.83% $\uparrow$ 0.008
C2 CNN $\uparrow$ 1.28% $\downarrow$ 0.017
C3 Hybrid $\uparrow$ 1.41% $\uparrow$ 0.022
Ours-SKM 2.91 84.07% 96.33% C1 CNN $\uparrow$ 3.73% $\uparrow$ 0.028
C2 CNN $\uparrow$ 2.57% $\uparrow$ 0.005
C3 Hybrid $\uparrow$ 2.70% $\uparrow$ 0.044
MRNet (Bien et al., 2018) Ours-S 8.46 73.88% 89.68% C1 CNN $\downarrow$ 0.52% $\downarrow$ 0.060
C2 CNN $\uparrow$ 0.00% $\downarrow$ 0.001
C3 Hybrid $\uparrow$ 0.53% $\uparrow$ 0.004
Ours-SK 9.20 76.81% 93.34% C1 CNN $\uparrow$ 2.65% $\uparrow$ 0.036
C2 CNN $\uparrow$ 1.58% $\uparrow$ 0.056
C3 Hybrid $\uparrow$ 2.12% $\uparrow$ 0.029
Ours-SKM 8.14 77.10% 93.81% C1 CNN $\uparrow$ 2.12% $\uparrow$ 0.030
C2 CNN $\uparrow$ 1.58% $\uparrow$ 0.009
C3 Hybrid $\uparrow$ 4.76% $\uparrow$ 0.035
4.8 Ablation Study
4.8.1 Analysis of Multimodal Conditions Guidance
To illustrate the role of different conditional guidance, we used five datasets to conduct comparative experiments with generators that resort to various banks of conditions for training and sampling. Table 6 validates that all proposed conditions are effective for enhancing synthesis and diagnosis tasks. Specifically, without text control, the downstream diagnostic performance is comparatively poor. Fig. 6(c-d) compares typical synthetic thyroid nodule sequences generated under condition banks with and without text guidance. The former is visibly high-fidelity and faithful to the given text prompt (e.g., smooth margin), while the latter exhibits less distinguishable features for diagnosis (e.g., blurry margin). That is to say, the proposed Ctrl-GenAug can create more diagnosis-reliable samples by enhancing semantic steerability in the generation process. Besides, by incorporating image prior knowledge in the sampling process, the three classifiers achieve consistent performance improvements on all datasets, with average accuracy/AUROC gains of Carotid (+4.40%, +0.045), TUSC (+4.20%, +0.040), ACDC (+1.87%, +0.019), MosMedData (+1.63%, +0.054), and MRNet (+2.64%, +0.017). This proves the domain gap between synthetic and real samples is mitigated by introducing the image prior. Moreover, as shown in Table 6, conditioning the generator on the motion field produces more content-consistent and smoother samples, leading to continuous improvements in diagnostic performance. It can also be observed that the FVD results show a limited correlation with downstream evaluation metrics, confirming the finding in Luo et al. (2024).
4.8.2 Impact of the Sequential Augmentation Module
To verify the impact of the proposed sequential augmentation module, as shown in Table 7, we compared the performance of three generator variants on the synthesis and downstream tasks. Ours-S, Ours-SK, and Ours-SKM denote our ablation studies, including gradually adding sequential attention (’-S’), key-frame/slice attention (’-K’), and motion field attention (’-M’) to the generator with spatial inflation only. Table 7 shows that the holistic downstream accuracy and AUROC improvements of Ours-SKM are significantly higher than those of Ours-S and Ours-SK in five datasets. In terms of quantitative assessment of sequential coherence, Ours-SKM achieves better results overall than other variants. These prove that this module can help the generator synthesize cross-frame/slice consistent and dynamic-smooth clips, which are diagnosis-promotive for downstream diagnostic tasks. Moreover, it can be observed that the FVD results across these three generators show no significant difference on four datasets, revealing its limited value for evaluating diagnosis-oriented medical sequence synthesis tasks.
Figure 8: Carotid stenosis and heart disease diagnostic accuracy (%) comparison using two classifiers on our filtering strategies under Real-finetune and Joint-train paradigms. The variant ’CF+SF’ indicates our final filter, and the accuracy gains over the Baseline are depicted.
4.8.3 Effectiveness of the Noisy Synthetic Data Filter